OpenStudy (anonymous):

f(x)= (x^2 + 1)^2 g(x)= (1)/(sqrt(x + 5) what is f o g what is g o f thanks in advance pls show all work

OpenStudy (anonymous):

\[f(x)= (x^2 + 1)^2\] \[g(x)= \frac{1}{\sqrt{x + 5}}\] \[f \cdot g=f(g(x))\]

OpenStudy (anonymous):

\[g \cdot f=g(f(x))=((\frac{1}{\sqrt{x + 5}})^2+1)^2\]

OpenStudy (anonymous):

You can do the second one.

OpenStudy (anonymous):

is the second one 1/(sqrt(x^3)^2 + 5)

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

Close, though.

OpenStudy (anonymous):

what did i do wrong then

OpenStudy (anonymous):

Sorry- my example was actually f(g(x))!

OpenStudy (anonymous):

Basically, to find the REAL g(f(x)) (not my mistaken one), write down g(x), and replace the x in g(x) with f(x).

OpenStudy (anonymous):